OpenStudy (anonymous):

identify a b c values y= 3/2x^2-x+8 y=7x-3x-1

OpenStudy (anonymous):

y = A x^2 + B x + C y = 3/2 x^2 -1 x + 8 A=3/2 B=-1 C=8 they are all like this... A is the coeficient of X^2 (the number before X^2) if there is no number, but the term is still there, then it is 1 if the term is missing, then it is 0

OpenStudy (anonymous):

a is the leading coefficient, the number in front of the highest power of x. In the first equation, a = 3/2. In the second, a = 4 because 7x-3x can be simplified to 4x. b is the coefficient in front of the second highest power of x. In the first equation, b = -1, since the x is negative. In the second equation, it is also -1, despite the number being a constant. c, as you might guess, is the coefficient of the next highest power of x after b. The first equation has a c of 8, the second equation has no c, or a c of 0.

OpenStudy (anonymous):

THANK YOU